37. Solve the quadratic equation by factoring.
x^2-9x-4=6

Answer by GANNU0716(2)


x^2-9x-4-6=0
x^2-9x-10=0
x^2-10x+x-10=0
x(x-10)+x-10=0
(x-10)(x+1)=0